OpenStudy (anonymous):

A cube has a surface area of 150 square yards. What is the length of a side? A)5 yd B)12.5 yd C)25 yd D)37.5 yd

OpenStudy (anonymous):

Formula for surface area of cube is 6a^2 where a = side length 6a^2 = 150 a^2 = 25 a = 5 yards

hartnn (hartnn):

surface area of cube = 6 * (side)^2 = 150 so, side =... ?

OpenStudy (anonymous):

5 yards

hartnn (hartnn):

yes.

OpenStudy (anonymous):

thank you guys for your help

OpenStudy (anonymous):

No Problem.